# A Review of Design Concerns in Superconducting Quantum Circuits

### v2

@@ V.2 Finite-element simulations
Figure 7: (a) Typical finite-element mesh that is seeded at the start of a simulation. Here the simulation is of a transmon with a central cross capacitor, an outer ground plane, and a “claw” coupling capacitor. The Josephson junction is modeled as a lumped element inductor (small green square at top of cross). During the solving, each mesh element is treated as having a single value of electric and magnetic field (or a spatial dependence given by a low-order polynomial), leading to a very inaccurate solution. (b) Mesh after several rounds of refinement. The refinement algorithm has correctly created a fine mesh on the transmon cross and coupling claw, but has not yet sufficiently refined the mesh on the ground plane. (c) Final mesh after solution convergence. The mesh elements are very small in any area where fields may be changing quickly.
@@ V.4 Simulation accuracy
Simulation convergence is determined by the change in some quantity (such as mode frequency) from one iteration to the next. A change in this quantity indicates that mesh refinement is significantly changing the solution, and that further refinement would likely further improve accuracy. Once the change between iterations is below some threshold, the convergence criteria are met and the simulation can end. The convergence threshold must be set to a low enough tolerance that the simulation accurately reproduces all quantities of interest, but high enough that the simulation finishes in a reasonable amount of time. Empirically we have found that a convergence threshold of less than .$05\%.$ change between iterations is good enough for most purposes [20].
@@
Unfortunately, most commercial solvers do not allow the user to pick which quantity is used to test convergence. This may prove to be an advantage for open-source solvers: advanced users can modify the back-end code and customize the convergence criteria. For example, a user may wish to tune the bandwidth $\kappa$ of a readout resonator to an accuracy within $5\%$. Ordinary eigenmodal simulations converge based on the changes in resonant frequencies $\Delta f$. In practice we have found that this level of $\kappa$ requires setting $\Delta f<0.05\%$ for robustness, but a coarser threshold may be sufficient in some situations. Unfortunately, it is not known if a coarser threshold will work before the simulation is done, and so best practice is to use a threshold that is always fine enough. A user could instead explicitly make convergence depend on $\kappa$, ensuring that the desired accuracy is achieved while minimizing computational cost.
@@
For complex layouts, it is generally better to simulate smaller sections instead of attempting to simulate the entire layout at once, which may be computationally prohibitive. These sections can then be turned into effective lumped circuit models [36] or effective quantized Hamiltonians and then combined. For example, for a design with a transmon capacitively coupled to a CPW cavity, one could simulate the transmon and coupling capacitor, then separately simulate the cavity and coupling capacitor. See Fig. 9 for an illustration. This modular approach allows for the extraction of accurate circuit parameters within each section and makes the simulations manageable in terms of both time and computational resources. The results from these smaller simulations can later be combined during the quantum analysis stage to construct the full system’s Hamiltonian. This method is particularly useful for large-scale designs with repeated subsystems, where simulating the entire structure at once would be impractical and computationally expensive to achieve the necessary level of accuracy. However, this approach may not capture all important effects in layouts where components are closely spaced. In such cases, cross-talk or hybridization of electromagnetic modes between neighboring elements may not be fully accounted for, potentially limiting the accuracy of the final analysis. Therefore, careful consideration should be given when applying this method to designs with tightly packed elements or strong mode interactions. The designer should make the simulation sections large enough that couplings outside the edges are negligible, or overlap them to ensure that boundary effects do not dominate.
@@
Figure 10: Demonstration of the dependence of simulation results on hyperparameters. Here we vary the number of consecutive passes that must satisfy the convergence threshold before the simulation can terminate. We compare the simulated resonant frequency (purple circles, left axis) and linewidth (orange squares, right axis) of a CPW resonator to the experimentally-measured values (dashed lines). In both cases the simulations saturate by 4 converged passes, but a small systematic error remains ($0.3\%$ for frequency and $8\%$ for linewidth) that can be attributed to fabrication and impedance variation effects.
@@ V.5.1 Energy participation ratio analysis
$\hat{H}_{\mathrm{nl}}$ $=\sum_{n=1}^{N}-E_{n}^{(J)}\left(\cos\hat{\phi}_{n}+\frac{\hat{\phi}_{n}^{2}}{2}\right)$ (17)
@@
The EPR method’s main advantages are its ease of use and its generality. It directly translates from eigenmodal simulation results to an effective diagonalized Hamiltonian with mode frequencies, self-Kerr terms, and cross-Kerr terms (similar to Section II). There is thus a direct translation between a physical layout and the observed spectrum. The pyEPR package [114] provides a simple Python API for running these calculations, and can be easily used by a designer who has no expertise in the physics behind the EPR method. EPR can capture an entire system’s behavior at once, reducing the number of required simulations. It makes few assumptions and adapts well to complex circuits with multiple nonlinear elements, making it a versatile tool in quantum circuit design. EPR can also be used to compute the loss caused by dissipative elements. However, it can be computationally demanding for large systems, as it requires simulation of a subsystem large enough that all relevant junctions are included for each mode. Additionally, there is some evidence that EPR analysis can be more vulnerable to inaccurate simulation [112, 20]. This is likely because EPR depends on accurate simulation of the field strength in each Josephson element, typically the smallest features in the design. The simulation may finish based on the solver’s convergence criteria while the junction fields are not yet accurately modeled. Modifying the solver to explicitly consider junction field in its convergence criteria may solve this issue (see Section V.4 for more discussion).
@@ VII Closing the design loop with measurements
One device properties have been characterized, it is important to determine whether any discrepancies with the target behavior are due to device design, fabrication, or embedding. Coherence measurements of control devices with standard designs can determine whether design is to blame for reduced coherence or whether it can be attributed to fabrication issues and/or improper filtering of measurement lines and/or improper shielding. Optical measurements can determine if the device geometry was fabricated as expected. Room-temperature resistance measurements of test structures can determine if Josephson energies were as designed. Once fabrication issues are eliminated, incorrect mode frequencies and couplings, excess crosstalk, and unwanted resonances can be confidently attributed to design. The most common reason for these missed parameters is some kind of issue with the simulation of the device—an electromagnetic solution had not converged to sufficient accuracy, or it was not interpreted as a Hamiltonian correctly. Using a robust and validated simulation pipeline is thus crucial for any device designer.
@@ VIII Outlook
Each topic covered in this short review is worthy of far more discussion, and the papers referenced here cover them in great detail. It is our hope that, until extensive didactic materials are developed, this review will function as a quick introduction for new designers and a central reference for experienced researchers looking to dig more into one topic.

### v3

@@ V.2 Finite-element simulations
Figure 7: (a) Typical finite-element mesh that is seeded at the start of a simulation. Here the simulation is of a transmon with a central cross capacitor, an outer ground plane, and a “claw” coupling capacitor. The Josephson junction is modeled as a lumped element inductor (small green square at top of cross). During the solving, each mesh element is treated as having a single value of electric and magnetic field (or a spatial dependence given by a low-order polynomial). When a mesh element is too large, this fails to capture significant field change within the element, leading to an inaccurate solution. (b) Mesh after several rounds of refinement. The refinement algorithm has correctly created a fine mesh on the transmon cross and coupling claw, but has not yet sufficiently refined the mesh on the ground plane. (c) Final mesh after solution convergence. The mesh elements are very small in any area where fields may be changing quickly, ranging from edge lengths of $0.1-1\ \mu$m at the edges of the ground plane to below $0.1\ \mu$m inside the qubit capacitor and coupling claw to below $0.01\ \mu$m inside the junction.
@@ V.4 Simulation accuracy
Simulation convergence is determined by the change in some quantities (such as mode frequencies or capacitances) from one iteration to the next. A large change in these quantities indicates that mesh refinement is significantly changing the solution, and that further refinement would likely further improve accuracy. Once the change in, e.g., mode frequency between iterations is below some threshold, the convergence criteria are met and the simulation can end. The convergence threshold must be set to a low enough tolerance that the simulation accurately reproduces all quantities of interest, but high enough that the simulation finishes in a reasonable amount of time. Empirically we have found that a convergence threshold of less than .$05\%.$ mean absolute change in mode frequency or capacitances between iterations is good enough for most purposes [22], where the mean is taken across all mode frequencies (for eigenmodal simulations) or all capacitance matrix elements (for electrostatic simulations).
@@
Another approach to improving accuracy is to add convergence criteria, explicitly mandating that the change in one or more additional quantities is less than some threshold. For instance, as shown in Fig. 8, mode frequency tends to converge faster than bandwidth. Mandating convergence on both may ensure an accurate solution without requiring as many extra passes after convergence. To the best of our knowledge there are no reported systematic studies of the advantages and drawbacks of using different convergence criteria, and we invite the community to rectify this!
@@
For complex layouts, it is generally better to simulate smaller sections instead of attempting to simulate the entire layout at once, which may be computationally prohibitive. For instance, on an in-house desktop workstation with an AMD Ryzen Threadripper PRO 3955WX 16-core 3.90 GHz processor and 128 GB of RAM, an accurate eigenmodal simulation of a 2x2 mm area takes between minutes and hours, while a 10x10 mm area can take days. These smaller sections can then be turned into effective lumped circuit models [38] or effective quantized Hamiltonians and then combined. For example, for a design with a transmon capacitively coupled to a CPW cavity, one could simulate the transmon and coupling capacitor, then separately simulate the cavity and coupling capacitor. See Fig. 9 for an illustration. This modular approach allows for the extraction of accurate circuit parameters within each section and makes the simulations manageable in terms of both time and computational resources. The results from these smaller simulations can later be combined during the quantum analysis stage to construct the full system’s Hamiltonian. This method is particularly useful for large-scale designs with repeated subsystems, where simulating the entire structure at once would be impractical and computationally expensive to achieve the necessary level of accuracy. However, this approach may not capture all important effects in layouts where components are closely spaced. In such cases, cross-talk or hybridization of electromagnetic modes between neighboring elements may not be fully accounted for, potentially limiting the accuracy of the final analysis. Therefore, careful consideration should be given when applying this method to designs with tightly packed elements or strong mode interactions. The designer should make the simulation sections large enough that couplings outside the edges are negligible, or overlap them to ensure that boundary effects do not dominate.
@@
Figure 10: Demonstration of the dependence of simulation results on hyperparameters. Here we vary the number of consecutive passes that must satisfy the convergence threshold before the simulation can terminate. We compare the simulated resonant frequency (purple circles, left axis) and linewidth (orange squares, right axis) of a CPW resonator to the experimentally-measured values (dashed lines). In both cases the simulations saturate by 4 converged passes, but a small systematic error remains ($0.3\%$ for frequency and $8\%$ for linewidth). The frequency error is mainly due to obfuscation of the fabrication etch bias, which are is a confidential parameter of the SQUILL foundry that fabricated the device. Including the etch bias reduces the error below 0.1% (purple star marker). The linewidth error is improved below 3% (orange star marker) when etch bias is included; we attribute the remaining error to impedance variations in the readout line, which simulations do not capture.
@@ V.5.1 Energy participation ratio analysis
$\hat{H}_{\mathrm{nl}}$ $=\sum_{n=1}^{N}-\left(E_{n}^{(J)}\cos\hat{\phi}_{n}+\frac{\hat{\phi}_{n}^{2}}{2}\right)$ (17)
@@
The EPR method’s main advantages are its ease of use and its generality. It directly translates from eigenmodal simulation results to an effective diagonalized Hamiltonian with mode frequencies, self-Kerr terms, and cross-Kerr terms (similar to Section II). There is thus a direct translation between a physical layout and the observed spectrum. The pyEPR package [117] provides a simple Python API for running these calculations, and can be easily used by a designer who has no expertise in the physics behind the EPR method. EPR can capture an entire system’s behavior at once, reducing the number of required simulations. It makes few assumptions and adapts well to complex circuits with multiple nonlinear elements, making it a versatile tool in quantum circuit design. EPR can also be used to compute the loss caused by dissipative elements. However, it can be computationally demanding for large systems, as it requires simulation of a subsystem large enough that all relevant junctions are included for each mode. Additionally, there is some evidence that EPR analysis can be more vulnerable to inaccurate simulation [115, 22]. This is likely because EPR depends on accurate simulation of the field strength in each Josephson element, which are typically the smallest features in the design. The simulation may finish based on the solver’s convergence criteria while the junction fields are not yet accurately modeled. Modifying the solver to explicitly consider junction field in its convergence criteria may solve this issue (see Section V.4 for more discussion).
@@ VII Closing the design loop with measurements
One device properties have been characterized, it is important to determine whether any discrepancies with the target behavior are due to device design, fabrication, or embedding. Coherence measurements of control devices with standard designs can determine whether design is to blame for reduced coherence or whether it can be attributed to fabrication issues and/or improper filtering of measurement lines and/or improper shielding. Optical measurements can determine if the device geometry was fabricated as expected. Room-temperature resistance measurements of test structures can determine if Josephson energies were as designed. Note that junction behavior can include departure from the single cosinusoidal energy expected for an ideal tunnel junction, which may alter spectroscopic results but not room-temperature resistance. Such behavior must be observed with low-temperature measurements [127].
Once fabrication issues are eliminated, incorrect mode frequencies and couplings, excess crosstalk, and unwanted resonances can be confidently attributed to design. The most common reason for these missed parameters is an inaccurate electromagnetic simulation of the device and/or an inaccurate translation of simulation results to Hamiltonian. Inaccurate simulation can be caused by, e.g., a simulation that does not include the whole geometry and thus fails to capture some long-range coupling or distributed mode; a simulation that does not have a tight enough tolerance for convergence or a fine enough mesh and thus produces an inaccurate solution; a simulation that fails to take into account the real geometry implemented including fabrication biases; or a simulation that does not take into account significant kinetic inductance. Inaccurate interpretation can be caused by, e.g. an interpretation that fails to account for finite-size effects of “lumped” elements that are not small compared to the wavelength; an interpretation that improperly treats distributed elements; an interpretation that truncates the basis space at too small of a dimension; an interpretation that uses approximate analytical formulas beyond their regime of accuracy; or an interpretation that fails to account for quantum effects such as phase slips. Using a robust and validated simulation pipeline is thus crucial for any device designer. Again the usual approach is empirical: if experimental results disagree with simulation, designers run larger simulations with tighter tolerances and add complexity to their interpretation models until agreement is reached.
@@ VIII Outlook
Each topic covered in this short review is worthy of far more discussion, and the papers referenced here cover them in great detail. In LABEL:tab:references we collect some of the key references, organized by topic. It is our hope that, until extensive didactic materials are developed, this review will function as a quick introduction for new designers and a central index for experienced researchers looking to dig more into one topic.
@@ Acknowledgements.
Table 1: References ordered by topic, with topics listed in order of appearance in the text. Some references unrelated to designs are omitted. A * next to an author name indicates a review article.
Topic Lead Author Year Reference
General superconducting qubits reviews Clarke* 2008 [2]
Devoret* 2013 [3]
Wendin* 2017 [4]
Krantz* 2019 [5]
Kjaergaard* 2020 [1]
Kwon* 2021 [6]
Blais* 2021 [7]
Gao* 2021 [8]
Rasmussen* 2021 [9]
Open-source codes Malinen 2013 [104]
Grozkowski 2021 [32]
Minev 2021 [119]
Minev 2021 [117]
Cucurachi 2021 [120]
Aumann 2022 [34]
Rajabzadeh 2023 [33]
Choi 2023 [105]
Shanto 2024 [22]
Shammah* 2024 [103]
Circuit quantization and diagonalization Ulrich 2016 [31]
Vool* 2017 [29]
You 2019 [10]
Kerman 2020 [30]
Chitta 2022 [35]
Riwar 2022 [11]
Bryon 2023 [12]
Rajabzadeh 2023 [33]
Weissler 2024 [13]
Cavity qubits reviews Joshi* 2021 [25]
Ma* 2021 [26]
Krasnok* 2024 [27]
Constructing circuit graphs and effective Hamiltonians from simulation Solgun 2014 [116]
Besedin 2018 [37]
Minev 2021 [28]
Minev 2021 [38]
Yuan 2022 [115]
Effects not captured by the rotating wave approximation Zueco 2009 [21]
Sank 2016 [19]
Dumas 2024 [20]
Scheuer 2014 [16]
Willsch 2017 [17]
Ann 2021 [18]
Classical crosstalk Abrams 2019 [39]
Dai 2021 [40]
Winick 2021 [44]
Wang 2022 [41]
Ketterer 2023 [42]
Yan 2023 [126]
Yang 2024 [43]
Quantum crosstalk Sheldon 2016 [55]
Yan 2018 [59]
Mundala 2019 [60]
Mitchell 2021 [56]
Kandala 2021 [61]
Wei 2022 [57]
Tripathi 2022 [58]
Air-bridge crossovers Abuwasib 2013 [45]
Chen 2014 [46]
Dunsworth 2018 [47]
Sun 2022 [48]
Janzen 2022 [49]
Tao 2024 [50]
Bu 2024 [51]
Flip-chip and 3D integration Rosenberg 2017 [52]
Foxen 2017 [62]
Vahidpour 2017 [63]
Yost 2020 [64]
Rosenberg* 2020 [65]
Li 2021 [53]
Kosen 2024 [54]
Connectivity and error correction Gottesman 2014 [71]
Short-range coupling schemes Moskalenko 2022 [67]
Weiss 2022 [69]
Ding 2023 [68]
Zhang 2024 [70]
Long-range coupling schemes Roch 2014 [77]
Schwartz 2016 [78]
Kurpiers 2018 [72]
Leung 2019 [73]
Kannan 2023 [74]
Zhou 2023 [76]
Greenfield 2024 [79]
Almanakly 2025 [75]
Dielectric loss Gao 2008 [80]
Paik 2011 [84]
Oliver* 2013 [81]
Wang 2015 [83]
Siddiqi* 2021 [82]
Martinis 2022 [85]
Smirnov 2024 [86]
Radiation and quasiparticles Barends 2011 [93]
Hays 2018 [90]
Houzet 2019 [89]
Rafferty 2021 [87]
Farmer 2021 [91]
Iaia 2022 [96]
Elfeky 2023 [92]
Liu 2024 [88]
Packaging Huang 2021 [95]
Signal injection in 3D cavities Axline 2016 [97]
Reshitnyk 2016 [99]
Stammeier 2018 [100]
Gargiulo 2021 [98]
Electromagnetic simulation Zmuidzinas* 2012 [110]
Kurra 2014 [108]
Naaman 2024 [106]
Hauer 2024 [111]
Pham 2023 [113]
Pham 2024 [112]
Characterization measurements Fink 2010 [125]
Willsch 2024 [127]